Can you help me with this difficult geometry problem?:
- BD bisects ∠ ABC. M∠ABC=7x. M∠ ABD=3x+25. Find M∠ DBC.

Answer by RAY100(1637)

m angle ABC = 7x,,,,,,bisector = 1/2 * 7x = 3.5 x
.
But bisector is angle ABD which is (3x+25)
.
setting these equal,,,,,3.5x = 3x+25,,,,,.5x=25,,,,,x=50
.
angle ABD = (3x+35) = 175 degrees,,,,,and bisector = 3.5(50) = 175 degrees
.
.
Angle DBC is also a bisector,,,therefore it also is 175 degrees
.